The decomposition reaction of NOBr is second order in NOBr, with a rate constant at 20°C of 25 M-1 min-1. If the initial concentration of NOBr is 0.025 M, find

(b) the concentration after 473 min of reaction.

M
1
Expert's answer
2016-11-03T08:14:54-0400
NOBr decomposition follows the equation 2NOBr = 2NO + Br2. Thus the rate equation for the second order reaction will be k2 = 1/t(1/C - 1/C0). Using this equation we can calculate the concentration after 473 min of the reaction:

C = C0/(C0*k2*t + 1) = 0.025 M/(0.025 M*25 M-1 min-1*473 min + 1) = 8.4*10-5 M.